The Hypotenuse of a right angled triangle is 25cm long if the other two sides are such that the length of one side is 5cm more than the length of one side find the length of the two sides
Found 2 solutions by JulietG, ikleyn:
Answer by JulietG(1812)

a^2 + b^2 = c^2 where c is the hypotenuse
a = b+5
Now your equation is
(b+5)^2 + b^2 = 625
b^2 +10b +25 +b^2 = 625
2b^2 + 10b +25 = 625
2b^2 +10b -600= 0
b^2 +5b -300 = 0
b = -15,20
Obviously there's no negative in measurement so ignore the sign. The sides are 15 and 20
15^2 + 20^2 = 25^2
225 + 400 = 625

Answer by ikleyn(52781)

.
The Hypotenuse of a right angled triangle is 25cm long if the other two sides are such that the length of one side
is 5cm more than the length of one side find the length of the two sides
---------------------------------------------------------

See the lesson Solved problems on the perimeter and side lengths of a right-angled triangle, Problem 4 in this site.